# Comparison of generalised additive models and neural networks in applications: A systematic review

### v1

@@ 7.1.3 AUC metric analysis
The marginal effects plot for different sample sizes through time over the empirical distribution is seen in Figure 7. The confidence bands for medium and unknown sample sizes include zero throughout, indicating no consistent difference in performance between GAMs and neural networks. In small datasets, GAMs initially outperform neural networks; however, this gap narrows through time. Conversely, neural networks had superior performance over GAMs for datasets with a large sample size, and this gap is narrowing over time. Note that neither the small nor the large datasets are statistically significant in recent years.
@@
The interaction between the number of predictors and year on the log AUC loss ratio is shown in Figure 8. Predicted marginal effects indicate minimal temporal change across different numbers of predictors. Datasets with a medium number of predictors show the most notable change through time. GAMs outperform neural networks in earlier years, with the gap narrowing to equivalence in 2024.
Although our analysis reveals some signals for variables impacting comparative model performance, the big variability in results across different papers is particularly noteworthy; as with the RMSE analysis, the ICC is large here (ICC =0.97=0.97). Thus, results on datasets from the same paper are highly correlated with each other, indicating differences in the modelling approaches and datasets specific to individual papers that have not been quantified.
@@ 7.1.4 $R^{2}$ metric analysis
Datasets utilising $R^{2}$ as a performance metric totalled 141, originating from 55 papers. The log $R^{2}$ loss ratio was modelled in the same manner as the above analyses. The inclusion of predictor variables and interactions provided no evidence of driving factors on the $R^{2}$ performance as seen in Table 9.
@@ 8 Conclusion
This systematic review and performance metric analysis evaluated 430 datasets across 143 papers comparing GAMs and neural networks, offering insights into how these two modelling frameworks have been applied and assessed in the literature. The first notable findings are the key characteristics of papers carrying out these comparisons and the substantial differences in how they are reported. Although all datasets included empirical analysis of both GAMs and neural networks, important dataset-level information, such as sample size, number of predictors, and neural network complexity, was frequently unreported. Neural network complexity, in particular, was only fully reported in 35% of datasets. The lack of standardised reporting affects transparency and reproducibility.
@@
We extracted and analysed performance outcomes across the most commonly reported metrics: RMSE, $R^{2}$, and AUC. On average, there was no evidence of a statistically significant difference between the performance of GAMs and neural networks across datasets for all three metrics based on the values of the intercepts in the models; that is to say, there is no evidence of consistent superiority of one model class over the other. This aligns with the claims of other authors, for example, Rudin, (2019) says that no single algorithm consistently outperforms others on tabular data, with differences often being small compared to tuning effects. They suggest interpretable models can perform as well as more complex ones, though publication bias often exaggerates performance gaps.
Considering covariate effects revealed potential drivers for differences in the performance of GAMs and neural networks. This highlights that, depending on the dataset structure and computing capabilities, one model may be preferred over the other on a case-by-case basis. For RMSE, model complexity and its interaction with year played a significant role. In earlier years, medium-sized neural networks were outperformed by GAMs, while neural networks of unknown size had superior performance over GAMs. Interestingly, while neural network size showed a significant association with the response, in recent years these effects have diminished and are no longer statistically distinguishable from zero for any network size.
For AUC, the pairwise interactions between year with sample size, and year with number of predictors showed that, while neural networks have tended to outperform GAMs in larger higher-dimensional datasets, this performance gap has narrowed over time; recent years show no statistically significant differences from zero for any category. This suggests that GAMs are highly competitive, especially in smaller data settings.
Our findings offer important insights for future research. First, GAMs remain a robust and interpretable choice in many applied settings and should not be dismissed simply due to the perceived superiority of more complex models. Second, in earlier years, while the performance advantage of models appeared dependent on neural network complexity and dataset dimension, in recent years, these factors no longer have an effect on performance. Researchers can weigh up the trade-off between model accuracy and interpretability. Third, the strong intra-paper correlation (high ICC values across all metrics) demonstrates the necessity of accounting for paper-level variation when comparing models. Given the large variability in model performance across different papers, it is of interest to investigate further potentially unrecorded factors at play.
@@
In conclusion, our study provides a comprehensive assessment of the performance of GAMs versus neural networks across a diverse set of applications using tabular data. While neural networks have shown impressive performance on complex (non-tabular) datasets, GAMs continue to offer competitive performance with added interpretability and transparency on tabular datasets. Neural networks and GAMs should be considered complementary tools within the modern data-science toolkit. Choosing between them requires careful consideration of the value placed on interpretability versus predictive gains. Notably, recent research has been conducted on combining GAMs and neural networks within hybrid models (Yang et al.,, 2021; Rügamer et al.,, 2024).
@@ Appendix A Alternative model formulation
In Section 7.1.2, the log RMSE ratio (GAM/NN) is modelled as
@@
No_Predictorsmedium  0.26 (0.22) -0.08 (0.24)  0.09 (0.37)
No_Predictorslarge  0.37 (0.34)  2.25 (0.72)  0.08 (0.72)
No_Predictorsunknown  0.10 (0.38)  0.88 (1.06) -0.20 (0.68)
@@
Year:Sample_Sizemedium -0.02 (0.03) -0.32 (0.02) -0.01 (0.05)
Year:Sample_Sizelarge -0.06 (0.04) -0.30 (0.05) -0.04 (0.10)
Year:Sample_Sizeunknown -0.10 (0.03) -0.21 (0.09) -0.01 (0.07)
Year:No_Predictorsmedium -0.03 (0.02) -0.03 (0.06) -0.03 (0.05)
Year:No_Predictorslarge -0.04 (0.04) -0.25 (0.09)  0.00 (0.07)
Year:No_Predictorsunknown -0.01 (0.04) -0.16 (0.13)  0.06 (0.07)
Year:Complexitymedium  0.11 (0.04)  0.05 (0.06)
Year:Complexitylarge  0.10 (0.05)  0.01 (0.07)
Year:Complexityunknown  0.01 (0.03) -0.06 (0.13) -0.03 (0.05)

### v2

@@ 7.4 AUC metric analysis
The marginal effects plot for different sample sizes through time over the empirical distribution is seen in Figure 8. The confidence bands for medium and unknown sample sizes include zero throughout, indicating no association with performance differences between GAMs and neural networks. In studies with small datasets prior to 2017, GAMs outperform neural networks, after which the difference is non-significant. Conversely, in studies with large datasets, there is a slight tendency for neural networks to outperform GAMs, with the effect becoming non-significant by 2015. Thus, in recent years, there is no association between sample size and model performance. This could be due to advancements in optimisation/tuning procedures in both model classes over time, but the specifics of the procedures used are not available as previously discussed.
@@
The interaction between the number of predictors and year also has a statistically significant effect on the log AUC loss ratio. The associated marginal effects are shown in Figure 9, where it is clear that the effect is relatively minimal. Datasets with a medium number of predictors show the most notable change through time, slightly favouring GAMs over neural networks, but with no difference in recent years.
Although our analysis provides some interesting signals that dataset dimensions (sample size and number of predictors) are associated with differences in model performance, the big variability in results across different papers is particularly noteworthy. As with the RMSE analysis, the ICC is large here (ICC =0.97=0.97). Thus, results on datasets from the same paper are highly correlated with each other, which may indicate differences in the modelling approaches and datasets specific to individual papers that have not been quantified.
@@ 7.5 $R^{2}$ metric analysis
Datasets utilising $R^{2}$ as a performance metric totalled 141, originating from 55 papers. The log $R^{2}$ loss ratio was modelled in the same manner as the RMSE and AUC analyses, with the results provided in Table 10. In the case of the $R^{2}$ metric, there were no statistically significant effects, but the ICC was large ($0.78$) in line with the other two metrics.
7.6 Sensitivity analysis
A series of sensitivity analyses were conducted to examine the stability of our performance-metric modelling under alternative data treatments and model specifications. We conducted a complete-case analysis, where cases with missing values in the sample size and number of predictors were excluded, and an analysis where these missing values were imputed using multivariate imputation by chained equations via the mice package (mice_package) in R. Note that, for both of these analyses, the neural network complexity variable was omitted due to its excessive amount of missingness. These analyses provide an alternative to the approach of including an “Unknown” category for missing values. We also carried out sub-analyses on the data from each of the four application domain areas separately, as an alternative to our previous overarching model wherein the domain appears as an additive effect. For these sub-analyses, we denote the fields as EAS (Environmental and Agricultural Sciences), EMCS (Engineering, Mathematics and Computer Science), HBS (Health and Biomedical Sciences), and OTH (Other). Across all of these variations, we find that the overall patterns in the results remain broadly consistent.
Table 11 displays the ICCs from all modelling approaches applied to all three metrics, where “Main” is the earlier model from Table 10. The ICC values remain high across the majority of scenarios, with values above 0.6 in almost all models. Interestingly, the lower ICC values relate to RMSE models, especially the complete data and HBS models. The generally large ICC values indicate significant differences between the results in different papers that cannot be explained by the available dataset-level variables. Moreover, the consistency in these high ICCs suggest that the clustering structure is an inherent feature of the literature (i.e., performance results are correlated within papers) rather than an artefact of any particular meta analysis model.
Table 11: ICC
Main Complete Impute EAS EMCS HBS OTH
RMSE $0.76$ $0.27$ $0.99$ $0.97$ $0.64$ $0.21$ $0.49$
AUC $0.97$ $0.97$ $0.74$ $0.76$ $0.88$ $0.93$ $0.94$
$\mathbf{R^{2}}$ $0.78$ $0.60$ $0.77$ $0.64$ $0.62$ $0.95$
Table 12: Likelihood ratio test p-values for each predictor across different meta analysis models for the three performance metrics.
Main Complete Impute EAS EMCS HBS OTH
Year RMSE 0.55 0.53 0.14 0.73 0.95 0.20 0.22
AUC 0.76 0.62 0.79 0.56 0.20 0.48 0.56
$R^{2}$ 0.39 0.59 0.61 0.26 0.90 0.97
Area RMSE 0.39 0.91 0.87
AUC 0.10 0.65 0.50
$R^{2}$ 0.71 0.89 0.90
Sample_Size RMSE $\mathbf{<0.01}$ 0.50 0.75 0.27 0.40 0.68
AUC $\mathbf{<0.01}$ $\mathbf{<0.01}$ 0.14 0.02 0.84
$R^{2}$ 0.83 0.73 0.88 0.55 0.64 0.81
Predictors RMSE 0.57 0.90 0.64 0.93 0.35 0.95
AUC 0.02 0.54 0.02 0.11
$R^{2}$ 0.84 0.31 0.63 0.30 0.45 0.95
Complexity RMSE $\mathbf{<0.01}$ $\mathbf{<0.01}$ 0.96 0.07
AUC 0.31 0.12
$R^{2}$ 0.60 0.06 0.20 0.48
Year $\times$ Area RMSE 0.69 0.92 0.38
AUC 0.73 0.17 0.43
$R^{2}$ 0.32 0.64 0.43
Year $\times$ Sample_Size RMSE 0.03 0.51 0.85 0.30 0.45
AUC $\mathbf{<0.01}$ 0.10 0.03 0.03 0.52
$R^{2}$ 0.92 0.70 0.86 0.51
Year $\times$ Predictors RMSE 0.69 0.71 0.66 0.83 0.31
AUC 0.01 0.17 0.06 0.13
$R^{2}$ 0.81 0.63 0.70 0.12
Year $\times$ Complexity RMSE 0.03 $\mathbf{<0.01}$ 0.87 0.09
AUC 0.65 0.09
$R^{2}$ 0.18 0.08
  • P-values less than 0.05 are emboldened for clarity, and effects that were inestimable in any analysis approach are indicated using the “—” symbol.
Significance of fixed effects was assessed using likelihood ratio tests for each of the seven analysis approaches, with resulting p-values provided in Table 12 (with the tables of regression coefficients deferred to B). The sample size effect over time is most often significant in the AUC models, whereas the effect of the number of predictors over time is only significant in the Main AUC model (which was a weaker effect from Figure 9). The imputation-based analysis identifies the effect of the sample size over time and the main effect of predictors (but where the time interaction has a p-value of $0.06$) in the AUC model, while the complete-case analysis only identifies the main effect of sample size. The key finding from the previous Main RMSE model was that the neural network complexity effect was significant. As this variable was removed from the complete-case and imputation-based analyses, it cannot appear in those models. However, its effect is significant in the RMSE model for the EAS papers. The AUC model for the EAS papers also identifies the effect of sample size as per the other AUC models, but not the effect of predictors (which, again, was the weaker effect). None of the other fields identify significant effects, but recall from Table 6 that the number of papers in each of these is smaller than for EAS (see also Table 14 in B); thus, the power to detect effects will be lower in each of these fields when considered individually. As for the $R^{2}$ metric, no variables are statistically significant in any analysis, in line with the Main $R^{2}$ model from before. We also conducted a version of the complete case analysis with results weighted by the sample size, which did not yield further insights (see B). Overall, our sensitivity analysis demonstrates that the main findings are relatively robust across variations in data handling and model specification.
@@ 8 Conclusion
This is the first systematic review of papers that compare GAMs and neural networks on real tabular datasets. In total, our review includes results from 430 datasets across 143 different papers, where we have also conducted an analysis of the performance metrics within these papers. Our work offers new insights into how these two modelling frameworks have been applied, assessed, and reported on within the literature.
Summarising the papers by year of publication, journal title and ranking, application area, problem type (regression/classification), sample size, number of predictors, and neural network type and complexity provides answers to RQ1 on the characteristics of papers comparing GAMs and neural networks. We found that the majority of papers in this area were published in top-ranking journals (approx. 70% in quartile one). No individual journal appeared more than three times, indicating a diversity of sub-disciplines represented from four broad domain areas (Environmental & Agricultural Sciences; Engineering, Mathematics & Computer Science; Health & Biomedical Sciences; and Other.) There has also been a recent surge of papers that compare GAMs and neural networks, particularly in the last five years. These findings highlight the growing interest in comparing these models across the scientific community, and timeliness of the current review.
Our study was not specific to any particular neural network architecture based on our search string or inclusion criteria (see Section 4). However, the abundance of studies using MLPs (approx. 90%) was particularly noteworthy, and 70% of these were shallow architectures consisting of a single hidden layer. While alternative neural network structures have been shown to improve on MLPs on tabular datasets, e.g., TabNet and Transformers (arik2021tabnet), comparisons with GAMs are lacking, and would be an interesting area of future research. Other recent studies show that tree-based procedures (e.g., random forests and XGBoost) outperform neural networks on tabular datasets (borisov2022deep; shwartz2022tabular). On the other hand, kadra2021well indicate that well-regularised MLPs can outperform other neural network structures and tree-based methods, suggesting that MLPs are a competitive baseline if appropriately regularised. Thus, MLPs remain relevant and are clearly very frequently used across diverse areas of applied research, but further comparison studies including modern machine learning methods, MLPs, and GAMs are of interest.
A very notable finding from RQ1 on the characteristics of papers carrying out these comparisons is substantial heterogeneity in how datasets and models are reported. Important details such as sample size, number of predictors, and neural network complexity were frequently unreported, with missingness of approximately 20% for the former two variables, and 65% for the latter. Other details on data preprocessing steps, optimisation algorithms, and hyperparameter tuning were inconsistently reported to the extent that they could not be captured in this review. It should be highlighted that missing information is a common feature of systematic reviews due to differences that will naturally arise in how different authors report results (andaur2022completeness). However, the lack of standardised reporting within machine learning research has previously been identified due to its impact on transparency and reproducibility (mitchell2019model). These reporting issues have been confirmed within our systematic review across many diverse research articles; the unreported information may also partly explain the high intra-paper correlation observed in our mixed-effects models. A particularly interesting future direction would be gathering datasets identified within this review, and independently analysing them according to a standardised modelling protocol, e.g., with consistent preprocessing, optimisation, tuning, and evaluation procedures to remove the effects of the different approaches taken in different papers.
@@
We also extracted and analysed performance outcomes across the most commonly reported metrics (RMSE, $R^{2}$, and AUC) to investigate both RQ2 (on whether or not one model outperforms the other) and RQ3 (and whether or not performance differences can be explained using reported variables). On average, there was no evidence of a statistically significant difference between the performance of GAMs and neural networks across datasets for all three metrics based on the values of the intercepts of our meta analysis models (RQ2). That is to say, there is no evidence of consistent superiority of one model class over the other. This aligns with the claims of other authors, for example, rudin2019stop says that no single algorithm consistently outperforms others on tabular data, with differences often being small compared to tuning effects. Indeed, Rudin suggests that interpretable models can perform similarly to complex ones, though publication bias often exaggerates performance gaps.
Considering the effects of paper and dataset characteristics revealed some potential associations with differences in the performance of GAMs and neural networks (RQ3). For RMSE, model complexity and its interaction with year was statistically significant. GAMs performed better than neural networks in studies where the neural network was medium-sized (5 to 10 parameters), but this effect has diminished in recent years and is no longer statistically significant. For AUC, the effects of dataset dimensions (sample size and the number of predictors) showed that, while neural networks are associated with improved performance in larger higher-dimensional datasets, the performance gap has narrowed over time. In recent years, there were no statistically significant effects associated with performance differences between GAMs and neural networks. While being conscious of the unreported information as discussed above, these results indicate that GAMs and neural networks are currently well matched across many different analyses conducted in the literature.
Our findings offer important insights for future research. First, GAMs remain a robust and interpretable choice in many applied settings and should not be dismissed simply due to the perceived superiority of more complex models. Indeed, GAMs are well suited to scientific research, where explainability and transparency are paramount; note that explainable machine learning is now a large field (montavon2019explainable; roscher2020explainable; burkart2021survey). Second, in the early-to-mid-2000s, while the performance advantage of models was associated with neural network complexity and dataset dimension, these factors no longer explain performance differences. Thus, it would appear that modern GAMs and neural networks cannot be distinguished based on these features, but it is possible that other features of optimisation and hyperparameter tuning may separate the two. Thus, future research should focus on adjusting for these additional modelling choices when conducting comparisons rather than only using one specific implementation of each model. Third, the strong intra-paper correlation (high ICC values across all metrics) demonstrates the necessity of accounting for paper-level variation when comparing models. Given the large variability in model performance across different papers, and inconsistency in reporting, we suggest that more standardised reporting of results should become the norm. An example of such a reporting standard is the “model card” proposed in mitchell2019model. This is particularly important as such models are increasingly being deployed (and automated) in mainstream applications, where transparency is becoming a focus (e.g., the EU’s AI Act).
@@
In conclusion, our study provides a comprehensive assessment of the performance of GAMs versus neural networks across a diverse set of applications using tabular data. We have also made the data collected available at https://github.com/jessicadoohan/gam-vs-nn-review, so that others can expand on this study in the future. While neural networks have shown impressive performance on complex (non-tabular) datasets, GAMs continue to offer competitive performance with added interpretability and transparency on tabular datasets. Neural networks and GAMs should be considered complementary tools within the modern data-science toolkit. Choosing between them requires careful consideration of the value placed on interpretability versus predictive gains. Notably, recent research has been conducted on combining GAMs and neural networks within hybrid models (Yang2021; NAM_timeseries2023; deepregression_R_paper_2023; rugamer2024semi).
@@ Appendix A Alternative model formulation
In Section 7.3, the log RMSE ratio (GAM/NN) is modelled as
@@
Predictorsmedium 0.26 (0.22) -0.08 (0.24) 0.09 (0.37)
Predictorslarge 0.37 (0.34) 2.25 (0.72) 0.08 (0.72)
Predictorsunknown 0.10 (0.38) 0.88 (1.06) -0.20 (0.68)
@@
Year × Sample_Sizemedium -0.02 (0.03) -0.32 (0.02) -0.01 (0.05)
Year × Sample_Sizelarge -0.06 (0.04) -0.30 (0.05) -0.04 (0.10)
Year × Sample_Sizeunknown -0.10 (0.03) -0.21 (0.09) -0.01 (0.07)
Year × Predictorsmedium -0.03 (0.02) -0.03 (0.06) -0.03 (0.05)
Year × Predictorslarge -0.04 (0.04) -0.25 (0.09) 0.00 (0.07)
Year × Predictorsunknown -0.01 (0.04) -0.16 (0.13) 0.06 (0.07)
Year × Complexitymedium 0.11 (0.04) 0.05 (0.06)
Year × Complexitylarge 0.10 (0.05) 0.01 (0.07)
Year × Complexityunknown 0.01 (0.03) -0.06 (0.13) -0.03 (0.05)
@@
Appendix B Sensitivity analysis
This appendix presents additional sensitivity analyses to assess the robustness of the main findings across alternative model specifications, subsets of the data, and weighting by sample size. Table 14 summarises how papers and datasets are distributed across model configurations for each performance metric.
Table 14: Number of papers and datasets for each model configuration.
Main Complete Impute EAS EMCS HBS OTH
RMSE 78 (227) 57 (149) 78 (227) 34 (66) 26 (112) 4 (14) 14 (35)
AUC 37 (111) 27 (65) 37 (111) 17 (81) 5 (6) 9 (10) 6 (14)
$\mathbf{R^{2}}$ 55 (141) 40 (98) 55 (141) 30 (60) 11 (44) 1 (6) 12 (30)
  • Papers are shown first with datasets shown in parentheses.
Table 15 reports mixed-effects model estimates based on the complete-case dataset, where observations with unknown sample sizes and number of predictors are removed. Table 16 shows estimates based on imputed data, i.e., where the unknown sample sizes and number of predictors are imputed. In both of these modelling approaches, the neural network complexity variable was removed due to its very high level of missingness.
Tables 17, 18, 19, and 20 respectively show the estimated coefficients from models fitted to each of the four domain areas separately, i.e., Environmental and Agricultural Science (EAS), Engineering, Mathematics and Computer Science (EMCS), Health and Biomedical Science (HBS), and Other (OTH).
Lastly, Table 21 fits a model where the observations are weighted by the sample size rather than including sample size as a covariate. This gives results from applications to larger datasets in the literature a larger weight in the meta analysis model. Note that this is conducted within the complete case analysis, where sample size is a known numeric variable, and, hence, the complexity variable was not included.
In most cases, individual coefficients are not statistically significant within these models. This does not necessarily mean that the overall covariates effects are not important, but rather that the differences compared to the reference categories are not statistically significant; overall covariate p-values were provided in Table 12 in the main paper. Note also that the number of papers analysed is naturally smaller in the complete case analysis and the various subset analyses, such that power to detect effects will be reduced. Interestingly, the EAS subset analysis (the largest subset) is closest to the full analysis in the main paper, i.e., the complexity-by-year interaction is statistically significant in RMSE model, and the sample size effect is statistically significant in the AUC model (but not its interaction with year, or the effect of predictor variables). In the analysis weighted by sample size, only one coefficient within the predictors-by-year interaction in the AUC model is statistically significant.
Table 15: Estimates from mixed-effects models using the complete dataset.
RMSE AUC $\boldsymbol{R}^{\boldsymbol{2}}$
Term Estimate (SE) Estimate (SE) Estimate (SE)
(Intercept) 0.02-0.02 (0.19) $\phantom{-}0.85$ (0.85) 0.68-0.68 (0.86)
Year $\phantom{-}0.00$ (0.02) $\phantom{-}0.05$ (0.10) $\phantom{-}0.04$ (0.08)
[…]
Sample_Sizemedium $\phantom{-}0.30$ (0.26) 0.10-0.10 (0.53) 0.17-0.17 (0.47)
Sample_Sizelarge $\phantom{-}0.21$ (0.38) 0.83-0.83 (0.52) 1.65-1.65 (2.51)
Predictorsmedium $\phantom{-}0.05$ (0.20) 0.17-0.17 (0.16) $\phantom{-}0.65$ (0.46)
Predictorslarge 0.12-0.12 (0.38) $\phantom{-}0.04$ (0.50) $\phantom{-}1.99$ (2.51)
[…]
Year × Sample_Sizemedium 0.03-0.03 (0.03) 0.03-0.03 (0.04) $\phantom{-}0.03$ (0.05)
Year × Sample_Sizelarge 0.01-0.01 (0.04) $\phantom{-}0.02$ (0.05) $\phantom{-}0.16$ (0.22)
Year × Predictorsmedium 0.01-0.01 (0.02) 0.06-0.06 (0.04) 0.03-0.03 (0.05)
Year × Predictorslarge $\phantom{-}0.01$ (0.04) 0.11-0.11 (0.06) 0.17-0.17 (0.22)
  • () indicates statistical significance at the 5% level.
Table 16: Estimates from mixed-effects models pooled from imputed datasets.
RMSE AUC $\boldsymbol{R}^{\boldsymbol{2}}$
Term Estimate (SE) Estimate (SE) Estimate (SE)
(Intercept) 0.22-0.22 (1.19) 0.10-0.10 (0.93) 0.38-0.38 (0.95)
Year $\phantom{-}0.11$ (0.12) $\phantom{-}0.04$ (0.13) $\phantom{-}0.05$ (0.09)
[…]
Sample_Sizemedium $\phantom{-}0.03$ (0.38) $\phantom{-}0.30$ (0.27) $\phantom{-}0.07$ (0.38)
Sample_Sizelarge $\phantom{-}0.40$ (0.66) $\phantom{-}0.21$ (0.26) $\phantom{-}0.14$ (0.82)
Predictorsmedium 0.01-0.01 (0.32) 0.06-0.06 (0.39) $\phantom{-}0.31$ (0.39)
Predictorslarge $\phantom{-}0.15$ (0.56) $\phantom{-}0.39$ (1.14) $\phantom{-}0.35$ (0.69)
[…]
Year × Sample_Sizemedium $\phantom{-}0.01$ (0.04) 0.05-0.05 (0.03) $\phantom{-}0.01$ (0.04)
Year × Sample_Sizelarge 0.02-0.02 (0.06) 0.04-0.04 (0.04) $\phantom{-}0.00$ (0.07)
Year × Predictorsmedium $\phantom{-}0.00$ (0.03) 0.02-0.02 (0.06) 0.02-0.02 (0.04)
Year × Predictorslarge 0.02-0.02 (0.05) 0.06-0.06 (0.14) 0.04-0.04 (0.07)
  • () indicates statistical significance at the 5% level.
Table 17: Estimates from mixed-effects models for the EAS subset analysis.
RMSE AUC $\boldsymbol{R}^{\boldsymbol{2}}$
Term Estimate (SE) Estimate (SE) Estimate (SE)
(Intercept) $\phantom{-}1.69$ (0.64) 2.33-2.33 (0.53) $\phantom{-}0.14$ (0.40)
Year $\phantom{-}0.08$ (0.09) $\phantom{-}0.47$ (0.15) $\phantom{-}0.02$ (0.05)
Sample_Sizemedium $\phantom{-}2.88$ (1.83) $\phantom{-}5.59$ (1.79) $\phantom{-}0.81$ (0.56)
Sample_Sizelarge 0.84-0.84 (2.88) 0.08-0.08 (2.71) $\phantom{-}0.31$ (1.33)
Sample_Sizeunknown $\phantom{-}1.86$ (1.36) $\phantom{-}2.68$ (0.40) 0.06-0.06 (0.61)
Predictorsmedium 0.02-0.02 (0.28) 15.39-15.39 (4.80) $\phantom{-}0.16$ (0.33)
Predictorslarge $\phantom{-}1.04$ (1.61) 12.04-12.04 (5.90) $\phantom{-}1.30$ (0.80)
Predictorsunknown $\phantom{-}0.04$ (3.10) $\phantom{-}2.83$ (1.72) 0.24-0.24 (1.07)
Complexitymedium 4.55-4.55 (1.05) $\phantom{-}14.36$ (5.97) 1.03-1.03 (0.56)
Complexitylarge 4.54-4.54 (1.04) 0.51-0.51 (0.72)
Complexityunknown 1.61-1.61 (0.91) 3.23-3.23 (1.87) $\phantom{-}0.44$ (0.31)
Year × Sample_Sizemedium 0.59-0.59 (0.31) 0.47-0.47 (0.14) 0.09-0.09 (0.07)
Year × Sample_Sizelarge $\phantom{-}0.05$ (0.26) 1.30-1.30 (0.89) 0.01-0.01 (0.12)
Year × Sample_Sizeunknown 0.10-0.10 (0.15) 0.17-0.17 (0.04) $\phantom{-}0.04$ (0.07)
Year × Predictorsmedium 0.04-0.04 (0.08) $\phantom{-}1.55$ (0.49) 0.05-0.05 (0.04)
Year × Predictorslarge 0.15-0.15 (0.18) $\phantom{-}0.85$ (0.46) 0.16-0.16 (0.08)
Year × Predictorsunknown 0.15-0.15 (0.28) 0.32-0.32 (0.16) 0.03-0.03 (0.11)
Year × Complexitymedium $\phantom{-}0.14$ (0.16) 1.29-1.29 (0.42) $\phantom{-}0.08$ (0.06)
Year × Complexitylarge $\phantom{-}0.38$ (0.09) $\phantom{-}0.04$ (0.07)
Year × Complexityunknown 0.03-0.03 (0.10) 0.06-0.06 (0.03)
  • () indicates statistical significance at the 5% level.
Table 18: Estimates from mixed-effects models for the EMCS subset analysis.
RMSE AUC $\boldsymbol{R}^{\boldsymbol{2}}$
Term Estimate (SE) Estimate (SE) Estimate (SE)
(Intercept) 0.47-0.47 (0.67) $\phantom{-}0.26$ (0.25) $\phantom{-}0.49$ (2.67)
Year $\phantom{-}0.00$ (0.05) 0.04-0.04 (0.03) 0.01-0.01 (0.11)
Sample_Sizemedium $\phantom{-}1.21$ (0.70) 0.32-0.32 (2.14)
Sample_Sizelarge $\phantom{-}0.45$ (0.59) 1.67-1.67 (1.31)
Sample_Sizeunknown $\phantom{-}0.97$ (1.58) 0.47-0.47 (2.09)
Predictorsmedium $\phantom{-}0.26$ (0.58) 0.84-0.84 (1.61)
Predictorslarge $\phantom{-}0.60$ (0.92) 3.01-3.01 (1.95)
Predictorsunknown 0.94-0.94 (0.59) 0.20-0.20 (1.69)
Total_Paramsmedium 0.93-0.93 (3.27)
Total_Paramslarge 0.15-0.15 (1.83) $\phantom{-}4.19$ (1.92)
Total_Paramsunknown $\phantom{-}0.24$ (1.15) $\phantom{-}0.33$ (1.50)
Year × Sample_Sizemedium 0.10-0.10 (0.07)
Year × Sample_Sizelarge 0.02-0.02 (0.06)
Year × Sample_Sizeunknown 0.08-0.08 (0.14)
Year × Predictorsmedium 0.03-0.03 (0.06)
Year × Predictorslarge 0.05-0.05 (0.09)
Year × Predictorsunknown $\phantom{-}0.09$ (0.06)
Year × Total_Paramsmedium $\phantom{-}0.12$ (0.28)
Year × Total_Paramslarge $\phantom{-}0.04$ (0.13)
  • () indicates statistical significance at the 5% level.
Table 19: Estimates from mixed-effects models for the HBS subset analysis.
RMSE AUC $\boldsymbol{R}^{\boldsymbol{2}}$
Term Estimate (SE) Estimate (SE) Estimate (SE)
(Intercept) 0.06-0.06 (0.02) $\phantom{-}0.13$ (0.29)
Year $\phantom{-}0.00$ (0.00) $\phantom{-}0.02$ (0.03)
  • () indicates statistical significance at the 5% level.
Table 20: Estimates from mixed-effects models for the OTH subset analysis.
RMSE AUC $\boldsymbol{R}^{\boldsymbol{2}}$
Term Estimate (SE) Estimate (SE) Estimate (SE)
(Intercept) 2.43-2.43 (4.50) 0.18-0.18 (0.73) $\phantom{-}1.25$ (2.37)
Year $\phantom{-}0.25$ (0.43) $\phantom{-}0.01$ (0.04) 0.01-0.01 (0.28)
Sample_Sizemedium $\phantom{-}0.38$ (0.44) 1.56-1.56 (2.99) $\phantom{-}0.64$ (2.56)
Sample_Sizelarge $\phantom{-}0.06$ (0.45) 0.30-0.30 (0.65) 0.54-0.54 (2.32)
Sample_Sizeunknown $\phantom{-}0.04$ (0.28) $\phantom{-}0.84$ (1.08)
Predictorsmedium 0.12-0.12 (0.34)
Predictorslarge 0.15-0.15 (0.32) 0.31 (0.93)
Predictorsunknown 0.03-0.03 (0.23) 0.32 (0.98)
Total_Paramsmedium $\phantom{-}2.72$ (4.37) 0.90-0.90 (1.70)
Total_Paramslarge 0.02-0.02 (5.11) 0.59-0.59 (2.41)
Total_Paramsunknown $\phantom{-}2.12$ (4.55) 1.78-1.78 (1.11)
Year × Total_Paramsmedium 0.27-0.27 (0.43)
Year × Total_Paramslarge 0.01-0.01 (0.50)
Year × Total_Paramsunknown 0.21-0.21 (0.44)
Year × Sample_Sizemedium 0.56-0.56 (0.73)
  • () indicates statistical significance at the 5% level.
Table 21: Estimates from mixed-effects models weighted by sample size.
RMSE AUC $\mathbf{R}^{2}$
Term Estimate (SE) Estimate (SE) Estimate (SE)
(Intercept) 0.14 (0.26) 0.02-0.02 (0.44) 0.90-0.90 (0.95)
Year 0.03-0.03 (0.03) 0.07 (0.07) 0.07 (0.08)
[…]
$\sigma^{2}_{\text{paper}}$ 0.10 (0.00) 0.14 $(0.00)^{*}$ 0.43 (0.00)
$\sigma^{2}_{\text{residual}}$ 229.85 (0.00) 79.49 (0.00) 2.40 (0.00)
ICC 0.00 (0.00) 0.97 $(0.00)^{*}$ 0.15 (0.00)
  • (*) indicates statistical significance at the 5% level.